Write in slope-intercept form the equation of the line that passes through the given points.
(2,3) and (4,7)

Answer by queenofit(120)

first you must find the slope
y2-y1
------ = m =slope your points are 1=(2,3)and 2=(4,7) put these in the formula
x2-x1
7-3
--- = 4/2=2 your slope is 2
4-2
then you must use the point-slope equation y-y1=m(x-x1)
y-3=2(x-2)
y-2=2x-4
y=2x-2 this is y=mx+b form which is your slope intercept form